OpenStudy (anonymous):

Fool's problem of the day, (1) Find the periodicity of \( sec(p \theta ) \csc(q \theta ) \) where \(p\) and \(q\) are rationals?

OpenStudy (experimentx):

multiplication of two periodic functions ...

OpenStudy (anonymous):

Yes so ?

OpenStudy (anonymous):

Hello Eyad, I will like to request you to refrain from making such invidious accusations. Medals are the least of the reasons for my participation in this site.

OpenStudy (experimentx):

seems unsolvable right now ...

OpenStudy (anonymous):

3.41 x 1,3019234 divided by 12347 x 8 + your age - 1804351 x 235-135425 squaredx 10 to the fifth power + 1- 1243818723847

OpenStudy (anonymous):

\(2\pi\), if \(p=q\) then \(\large\frac{\pi}p\).